OpenStudy (anonymous):

What is the midpoint of the line segment whose endpoints are (-8, 12) and (-13, -2)? (-10.5, 5) (-10.5, 7) (-2.5, 7)

OpenStudy (amriju):

a

OpenStudy (solomonzelman):

\[\large\left(\begin{matrix} \\ \frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\end{matrix}\right)\]

OpenStudy (solomonzelman):

you don;t first give the answer and then explain. I'll report you next time !